OpenStudy (anonymous):

check please find the area of a triangle ABC if a=40 ft, b=25 ft, and angle C=30 degrees i said 250ft^2

OpenStudy (anonymous):

area of triangle is given by 1/2*a*b*sinC=1/2 *40*25*sin(30) =1/2 *40*25*1/2=250

OpenStudy (anonymous):

@nisa correct...

OpenStudy (anonymous):

thankssss :D

OpenStudy (anonymous):

mention not as u already got correct answer

OpenStudy (anonymous):

yes but u took ure time to see it and check it :)

OpenStudy (anonymous):

thats not big deal for maths lovers

OpenStudy (anonymous):

yes but for me it is im not too good in math and knowing i did something right gives me a thrill lol and that somebody took the time to check it :)